OpenStudy (anonymous):

Determine the inverse f^-1 of f when F(x)= 3+2x^3

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

solve for \(y\) in \(x=3+2y^3\)

OpenStudy (anonymous):

\[x=3+2y^3\\ x-3=y^3\\ \sqrt[3]{x-3}=y\]

OpenStudy (anonymous):

your answer is \[f^{-1}(x)=\sqrt[3]{x-3}\]

OpenStudy (anonymous):

ooh no i forgot about the two!!

OpenStudy (anonymous):

\[x=3+2y^3\\ x-3=2y^3\\ \frac{x-3}{2}=y^3\\ \sqrt[3]{\frac{x-3}{2}}=y\]

OpenStudy (anonymous):

that's better your answer is \[f^{-1}(x)=\sqrt[3]{\frac{x-3}{2}}\]

OpenStudy (anonymous):

Can you help me with hw?

OpenStudy (anonymous):

I can give you my login info

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

Can you help me with hw?

OpenStudy (anonymous):

I don't mind giving you my login info for the site I really need help

OpenStudy (anonymous):

The URL is there

OpenStudy (ybarrap):

Site says "Unauthorized use of UT Austin computer and networking resources is prohibited. If you log in, you acknowledge your awareness of and concurrence with the UT Austin Acceptable Use Policy. The university will prosecute violators to the full extent of the law. ..." I don't think it's wise to give out this info.

OpenStudy (anonymous):

Okay one moment

OpenStudy (anonymous):

@ybarrap can I get your email so I can send you the problem?

OpenStudy (anonymous):

@yb

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

@michelle010d - Just post your problems -- these aren't too difficult and you would be able to get a lot more help.